# AI for the Open-World: the Learning Principles

### v1

@@ 4.2.1 Memory Mosaics v2 architectures
Compared with the Memory Mosaics architecture in section 3.5, Memory Mosaics v2 incorporates three main modifications, including: 1) 3-level memories, 2) adaptive bandwidth, and 3) gated key feature extractor. This section provides a detailed explanation of the Memory Mosaics v2 architecture.
Figure 4.2: (left) Transformer architecture; (right) Memory Mosaics v2 architecture
4.2.1.1 3-level memories
Memory Mosaics v2 consists of three memories: 1) short-term memory, 2) long-term memory, 3) persistent memory, as shown in Figure 4.2. This hierarchical design distributes features to different memories according to their level of invariance. For example, a feature like “people can speak and eat” is more invariant than a feature like “I am reading a paper”, and therefore should be stored in persistent memory. The goal of this hierarchy feature allocation is to maximize feature reuse during new task learning. Ultimately, help a model learn new tasks with fewer examples and less priori knowledge from designers.
Persistent memory, which stores persistent knowledge, is implemented using a Feed-Forward Network due to computational efficiency concerns. [^7] A two-layers feed-forward network and a key-value associative memory are interchangeable as shown in Sukhbaatar et al. [2019]. The formula is given by:
$F_{p}(x_{t})=W_{3}^{\top}\Bigl{(}\text{SiLU}\left(W_{1}^{\top}\text{RMSNorm}(x_{t})\right)\odot W_{2}^{\top}\text{RMSNorm}(x_{t})\Bigr{)}\,,$ (4.1)
where $\text{SiLU}(x)=x\cdot\sigma(x)$ is sigmoid linear unit function, $x\in\mathbb{R}^{n}$, $W_{\{1,2\}}\in\mathbb{R}^{n\times m},\,W_{3}\in\mathbb{R}^{m\times n}$.
Long-term memory, which stores temporary working knowledge, is an associative memory that stores past key-value pairs $\{(k_{1},v_{1}),\dots,(k_{t-m},v_{t-m})\}$ starting from a delayed starting time point $t-m$ (Figure 4.3). The formula for long-term memory is given by:
$F_{l}(x_{t})$ $=W_{l}\,f\Bigl{(}k_{t};\tau_{m}\bigl{(}[(k_{1},v_{1}),\dots,(k_{t-1},v_{t-1})]\bigr{)}\Bigr{)}$ (4.2)
$\tau_{m}([x_{1},\dots,x_{t-1}])$ =$=\begin{cases}[x_{1},\dots,x_{t-m}]&\text{if\, $m>t$}\\ []&\text{otherwise}\end{cases}\,,$ (4.3)
where $f(\cdot)$ is a Guassian kernel smoothing function that will be introduced later, $\tau_{m}(\cdot)$ truncates the latest $m$ examples of a sequence, and $W_{l}$ is learnable parameter.
Figure 4.3: (left) Long term memory; (right) corresponding attention mask.
Short-term memory is also an associative memory, but only stores a recent past, i.e., key-value pairs $\{(k_{t-h},v_{t-h}),\dots,(k_{t-1},v_{t-1})\}$. The formula for short-term memory is given by:
$F_{s}(x_{t})$ $=W_{s}\,f\Bigl{(}k_{t};\kappa_{h}\bigl{(}[(k_{1},v_{1}),\dots,(k_{t-1},v_{t-1})]\bigr{)}\Bigr{)}$ (4.4)
$\kappa_{h}([x_{1},\dots,x_{t-1}])$ =$=\begin{cases}[x_{t-h},\dots,x_{t-1}]&\text{if $h<t$}\\ [x_{1},\dots,x_{t-1}]&\text{otherwise}\end{cases}$ (4.5)
where $\kappa_{h}(\cdot)$ keeps the latest $h$ examples of a sequence, and $W_{s}$ is learnable parameter.
Figure 4.4: (left) Short term memory; (right) corresponding attention mask.
@@ 4.2.1.2 Adaptive bandwidth in Gaussian kernel smoothing
In Memory Mosaics v2, the “memory” in long-term and short-term memory is the same as the associative-memory in Memory Mosaics in Section 3.2.1. The only difference is that Memory Mosaics v2 uses an adaptive bandwidth parameter $\beta$:
@@
$\beta$ $=\beta_{1}i^{\alpha}+\beta_{0}$ (4.7)
where $\|k_{i}\|=\|v_{i}\|=1$, $1/\sqrt{\beta}$ is propositional to bandwidth in Gaussian kernel smoothing. Inspired by the Asymptotic MISE and plug-in rule in kernel bandwidth estimation [García-Portugués 2024] ($1/\sqrt{\beta}\propto i^{-1/(p+4)}$), we scale $\beta$ as $\beta=\beta_{1}i^{\alpha}+\beta_{0}$, where $\beta_{0},\,\beta_{1},\,\alpha$ are learnable parameters.
4.2.1.3 Key and Value Feature Extractor
Value $v_{t}$ is constructed as convolution on time $t$ and $t+1$. Inspired by rwkv and mamba [Peng et al. 2023b; Gu and Dao 2023b], Key $k_{t}$ at time $t$ are extracted through a gated RNN structure, as illustrated in Figure 4.5. The formula is give by:
$k_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}\big{)}\quad\text{with}\quad\overbrace{\begin{cases}{\,{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}=g_{T}{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}+\lambda_{T}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}\\ \,g_{t}=e^{W_{g}x_{T}}\in\mathbb{R}\,,\,\lambda_{T}=e^{-|W_{\lambda}x_{T}|}\in\mathbb{R}\\ \end{cases}}^{\text{leaky average over t = T, {T-1}\ldots, 1}}$ (4.8)
@@ 4.2.1.3 Key and Value Feature Extractor
$v_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}\big{)}\quad\,\text{with}\quad\quad\underbrace{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}={\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}+\lambda_{\psi}{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T+1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}=W_{\psi}\,x_{T}}_{\text{convolution overt=T and T+1}}$
where $W_{\varphi},W_{g},W_{\lambda},W_{\psi},\lambda_{\psi}$ are learnable parameters.
Figure 4.5: Key $k_{t}$ feature Extractor: $k_{t}=\text{Norm}(\bar{k}_{t})$
Please note that we explicitly separate long-term memory and short-term memory without any shared parameters. This design not only helps the allocation of different features into different memories, but also provide engineering convenient in replacing the implementation of memories. For example, replacing short-term memory with rwkv or sliding-window attention, replacing long-term memory with hash-table or sparse attention.
@@ 4.2.2 Large-scale Training
We train two Memory Mosaics v2: Memory Mosaics v2 Small and Memory Mosaics v2 Large. Memory Mosaics v2 Small contains 24 layers, 2048 hidden dimensions, and 16 heads. Memory Mosaics v2 large increases the number of layers to 32, hidden dimensions to 4096, and the number of heads to 32.888The persistent memory dimension $m$ (in Equation 4.1) is set to 6,144 and 14,336 for Small and Large models, respectively.
@@
It is worth noting that these hyperparameters were originally searched and optimized for the baseline Transformer models (by amaia team). We transfer these hyperparameters to Memory Mosaics v2 without further hyperparameter searching. Thus, it is possible that this hyperparameter setup is suboptimal for Memory Mosaics v2.999A suboptimal hyperparameter on Memory Mosaics v2 is acceptable for this thesis, because Memory Mosaics v2 experiments are simply used to verify the inference-time learning principle.
@@ 4.2.3 Three evaluation dimensions
  • Persistent-knowledge storage and retrieval. This dimension evaluates the ability of persistent-memory to store and retrieve knowledge. This ability, storing and retrieval knowledge of training data, prepares features/knowledge that could be reused on other tasks during inference. We use common language model benchmarks to access this aspect.
  • New-knowledge storage and retrieval. This dimension assesses the ability to store and retrieve new information in inference-time. It is a prerequisite for “learning” new tasks via memory-based methods. We employ "needle-in-a-haystack" tasks to evaluate this aspect. [^11] Imagining a poor goldfish with a 7-seconds long-term memory, how can it learn a 90-mins movie? Various approaches with fix-sized memories, including RNNs, LSTM, state-stace model, sliding-window attention, resemble this poor goldfish. Memory Mosaics v2 stores new knowledge in long-term memory, while transformers store new knowledge in attention.
  • In-context Learning. This dimension directly evaluates the ability of the model to learn new tasks with fewer examples and less task-specific priori knowledge. We use multiclass classification to assess this aspect.
@@ 4.2.3.1 Persistent-knowledge storage and retrieval
This section first evaluates both Memory Mosaics v2 and baseline transformers on 19 commonly-used language benchmarks. Table 4.1 shows that Memory Mosaics v2 and transformers performs closely on these benchmarks. This is nothing to supervise, because both Memory Mosaics v2 and transformers use the same persistent memory architecture. Furthermore, many other architectures, such as mamba [Gu and Dao 2023b] and xlstm [Beck et al. 2025], also perform well on these benchmarks.
@@
Table 4.2 and 4.3 summarize the number of parameters and computation required for transformers and Memory Mosaics v2.131313The flops per token is estimated via Casson [2023] approach. Compared with Transformer Large, Memory Mosaics v2 Large uses slightly more parameters and computations to explicitly allocate 3-level memories (persistent-memory, long-term memory and short-term memory). This 3-levels design is used to allocate features according to their level of invariance, help feature reusing on new tasks. Removing long-term memory of Memory Mosaics v2 after training achieves a comparable performance of transformer on the 13 persistent-knowledge benchmarks, while using fewer parameters and computations.
@@ 4.2.3.2 New-knowledge storage and retrieval
The new-knowledge storage and retrieval ability is a prerequisite for learning new tasks via memory-based methods, because the data of new tasks should be well “stored” before learning (memory-based methods are lazy methods). To illustrate this point, consider a poor goldfish with 7-second memory — how can it learn a 90-minute movie? Similarly, a model with limited new-knowledge storage ability will struggle to learn from data that exceed its storage (memory) capacity. To assess this ability, we employ two “needle in a haystack” question-answering tasks from the ruler benchmark [Hsieh et al. 2024]. These tasks involve multiple concatenated (real) articles followed by a question related to one of the articles, requiring the model to find the correct answer based on the correct article.
@@
The two question-answer tasks [Hsieh et al. 2024] consist of multiple articles collected from SQuAD [Rajpurkar et al. 2016] and HotpotQA [Yang et al. 2018] (realistic articles), followed by a question corresponding to a certain article. The model must find the correct answer according to the relevant article. By concatenating more (unrelated) articles, the task-length becomes larger and the task itself becomes more difficult. An example prompt is shown below:
@@
Similarly to the process used in Section 4.2.3.1 for verifying persistent-knowledge storage and retrieval tasks, Appendix Table C.1 compares Memory Mosaics v2 with and without long-term memory on these question-answering tasks, verifying that these tasks truly require “long-term memory”.
Table 4.4 compares Transformer and Memory Mosaics v2, pretrained on 4k context length, evaluated on the question-answer tasks. Memory Mosaics v2 outperforms Transformer on 4k task-length by 1.4$\sim$5.6%. Similarly, Table 4.5 presents the same comparison but with both Transformer and Memory Mosaics v2 models fine-tuned on 32k context length. After context-length extension, Memory Mosaics v2 significantly outperforms Transformer on 32k task-length by 12.3$\sim$14.8%.

### v2

@@ 4.2.1 Memory Mosaics v2 architectures
Compared with the Memory Mosaics architecture in section 3.5, Memory Mosaics v2 incorporates three architecture modifications, including an adaptive bandwidth in associative memory, a gated time-variant key feature extractor, and a 3-level memory design. This section provides a detailed explanation of the Memory Mosaics v2 architecture.
@@ 4.2.1.1 Adaptive bandwidth in Gaussian kernel smoothing
Memory mosaics use one fixed bandwidth parameter $\beta$ for different sizes $n$ of associative memory (Equation 3.12). It is well known that bandwidth controls the bias-variance trade-off [Hastie et al. 2009] of kernel regression (memory-based) methods. That is, for a given distribution, the optimal bandwidth depends on the number of examples (key-value pairs in associative memory). Inspired by the asymptotic Mean Integrated Squared Error kernel bandwidth estimation approach where $1/\sqrt{\beta}\propto n^{-1/(p+4)}$ [García-Portugués 2024], memory mosaics v2 scale $\beta$ in associative memories as: [^7] For the easy of reading, this chapter reuse the “attention score” notion to $\frac{e^{\,\beta\,k^{\!\top}k_{i}}}{\sum_{j=1}^{n}e^{\,\beta\,k^{\!\top}k_{j}}}$ in associative memories.
@@
$\beta$ $=\beta_{1}n^{\alpha}+\beta_{0}$ (4.2)
[…]
4.2.1.2 Gated time-variant key feature extractor
Memory mosaics in Section 3.2.1 employ a simple time-invariant leaky averaging to extract key features:
@@ 4.2.1.2 Gated time-variant key feature extractor
$k_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}\big{)}$ $\text{with}\quad{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}={\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}+\lambda\,{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}$ (4.3)
The averaging weights $\lambda$ in Equation 4.3 are fixed and independent of the semantic input $x$. As a result, semantically similar cases, such as “tom-and-jerry” and “tom- - -and- - -jerry”, may receive different key features. Inspired by recurrent-style networks [Peng et al. 2023b; Gu and Dao 2023b; Beck et al. 2025], memory mosaics v2 utilize the following gated time-variant key feature extractor: [^8] It worth noting that this work is neither a linearization of attention nor attention efficiency. The recurrent feature extractor in Eq. 4.3 is used to create keys, while associative memory in Eq. 4.1 still stores all key-value pairs.
$k_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}\big{)}\quad\text{with}\quad\begin{cases}{\,{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}=g_{T}{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}+\lambda_{T}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}\\ \,g_{t}=e^{W_{g}x_{T}}\in\mathbb{R}\,,\,\lambda_{T}=e^{-|W_{\lambda}x_{T}|}\in\mathbb{R}\end{cases},$ (4.4)
where $W_{\varphi},W_{g},W_{\lambda}$ are learnable parameters, the averaging weights $\lambda_{T}\in\mathbb{R}$ and the the exponential gate $g_{T}\in\mathbb{R}$ semantically depend on input $x_{T}$.
For key feature extractor, memory mosaics v2 reuses the same convolutional key extractor as in memory mosaics:
$v_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}\big{)}\quad\,\text{with}\quad{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}=\gamma\,{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}+(1-\gamma)\,{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T+1}}\quad~{}{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}=W_{\psi}\,x_{T}\,,}$ (4.5)
where $\gamma\in\mathbb{R}$ and $W_{\psi}$ are learnable parameters.
4.2.1.3 3-level memory
The memory mosaics in Section 3.2.1 simplify the attention in transformer as contextual associative memory, view the feed-forward network in transformer as persistent memory. This simplification reduces the dependence between the “attention score” and the token position, as shown in Figure 4.2.
Figure 4.2: Average attention scores of the last token attending previous tokens. Left: Transformer with RoPE position encoding. Right: Memory Mosaics in Section 3.2.1. The (averaged) attention scores in transformer heavily depends on token positions (curly curves), while the attention scores in memory mosaics at far tokens (e.g. position 0 to 450) are almost invariant to positions (flat curves).
Compared with transformers (figure 4.2 left), the attention scores in memory mosaics exhibit a structured pattern (figure 4.2 right). That is, attention scores on near-tokens (positions) heavily depends on positions, while attention scores on far-tokens are almost invariant to token positions. Inspired by this experimental discovery, memory mosaics v2 replace contextual associative memory in memory mosaics with two associative memories, short-term memory and long-term memory, using distinct parameters (as figure 4.3). The short-term memory at position $t$ only store key-value pairs of near tokens, ranging from $t-h+1$ to $t-1$. In contrast, the long-term memory skips near tokens and only store key-value pairs before position $t-m$. By setting $m<h$, memory mosaics v2 create an overlap between long-term and short-term memory, resulting in a soft boundary between these two memories.
Memory mosaics v2 implements persistent memory using dense two-layers neural networks with SwiGLU activation [Shazeer 2020] due to computational efficiency concerns. [^9] A two-layers feed-forward network and a key-value associative memory are interchangeable as shown in Sukhbaatar et al. [2019].
Figure 4.3: Memory Mosaics v2 architecture.
@@ 4.2.2 Large-scale Training
We train two Memory Mosaics v2: Memory Mosaics v2 Small and Memory Mosaics v2 Large. Memory Mosaics v2 Small contains 24 layers, 2048 hidden dimensions, and 16 heads. Memory Mosaics v2 large increases the number of layers to 32, hidden dimensions to 4096, and the number of heads to 32.101010The persistent memory hidden dimension is set to 6,144 and 14,336 for Small and Large models, respectively.
@@
It is worth noting that these hyperparameters were originally searched and optimized for the baseline Transformer models (by amaia team). We transfer these hyperparameters to Memory Mosaics v2 without further hyperparameter searching. Thus, it is possible that this hyperparameter setup is suboptimal for Memory Mosaics v2.111111This hyperparameter setup is sufficient for memory mosaics v2 to verify the inference-time learning principle.
@@ 4.2.3 Three evaluation dimensions
  • Persistent-knowledge storage and retrieval, the ability of persistent-memory to store and retrieve knowledge of training dataset. This capability prepares knowledge that could be reused in other tasks during inference. We use common language benchmarks to access this aspect.
  • New-knowledge storage and retrieval, the ability to store and retrieve new information of test dataset. It is a prerequisite for “learning” new tasks via memory-based methods. We employ “multi-unrelated-documents storing and question-answering” tasks to evaluate this aspect. [^13] Imagining a poor goldfish with a 7-seconds long-term memory, how can it learn a 90-mins movie? Various approaches with fix-sized memories, including RNNs, LSTM, state-stace model, sliding-window attention, resemble this poor goldfish.
@@ 4.2.3.1 Persistent-knowledge storage and retrieval
This section first evaluates both Memory Mosaics v2 and baseline transformers on 19 commonly-used language benchmarks. Table 4.1 shows that Memory Mosaics v2 and transformers performs closely on these benchmarks. This is nothing to supervise, because both Memory Mosaics v2 and transformers use the same persistent memory architecture. Meanwhile, many other RNN architectures, such as mamba [Gu and Dao 2023b] and xlstm [Beck et al. 2025], also perform well on these benchmarks.
@@
Table 4.2 and 4.3 summarize the number of parameters and computation required for transformers and Memory Mosaics v2.151515The flops per token is estimated via Casson [2023] approach. Compared with Transformer Large, Memory Mosaics v2 Large uses slightly more parameters and computations to explicitly allocate 3-level memories (persistent-memory, long-term memory and short-term memory). It is worth paying more parameters in exchange for the 3-level memory design, because this design helps in allocating features according to their level of invariance, facilitating feature reuse in new tasks. This 3-levels design is used to allocate features according to their level of invariance, help feature reusing on new tasks. Furthermore, removing long-term memory from memory mosaics v2 after training achieves a comparable transformer performance on the 13 persistent-knowledge benchmarks, while using fewer parameters and computations.
@@ 4.2.3.2 New-knowledge storage and retrieval
The new-knowledge storage and retrieval ability is a prerequisite for learning new tasks via memory-based methods (e.g., Gaussian kernel regression), because the data of new tasks must be adequately “stored” before learning (Note that memory-based methods are lazy methods). To illustrate this point, consider a poor goldfish with 7-second memory – how can it possibly learn a 90-minute movie? Similarly, a model with limited new-knowledge storage ability will struggle to learn information that exceeds its storage (memory) capacity.
@@
To assess this ability, we employ two “multi-unrelated-documents question-answering” tasks from the ruler benchmark [Hsieh et al. 2024]. These tasks involve multiple concatenated realistic articles, collected from SQuAD [Rajpurkar et al. 2016] and HotpotQA [Yang et al. 2018], followed by a question related to one of these articles. Then expect the model to find the correct answer based on the correct article. [^16] Similarly to the process used in section 4.2.3.1 for verifying persistent-knowledge storage and retrieval tasks, appendix Table C.1 compares memory mosaics v2 with and without long-term memory on these question-answering tasks, confirming the necessity of “long-term memory” for these tasks. By concatenating more (unrelated) articles, the task-length becomes larger and the task itself becomes more challenging. An example prompt is shown below:
@@
Similarly to the process used in Section 4.2.3.1 for verifying persistent-knowledge storage and retrieval tasks, Appendix Table C.1 compares Memory Mosaics v2 with and without long-term memory on these multi-unrelated-documents question-answering tasks, verifying that these tasks truly require “long-term memory”.
Table 4.4 compares Transformer and Memory Mosaics v2, pretrained on 4k context length, evaluated on the question-answer tasks. Memory Mosaics v2 outperforms Transformer on 4k task-length by 1.4$\sim$5.6%. Similarly, Table 4.5 presents the same comparison but with both Transformer and Memory Mosaics v2 models fine-tuned on 32k context length. After context-length extension, Memory Mosaics v2 significantly outperforms Transformer on 32k task-length by 12.3$\sim$14.8%. Moreover, memory mosaics v2 also outperforms many other public base models of similar scale (See Appendix C.1 Table C.2 for the details).